Munoz Company started year 1 with $210,000 in its cash and common stock accounts. During year 1, Munoz paid $157,500 cash for employee compensation and $48,300 cash for materials.

Required:

A) Determine the total amount of assets and the amount of expense shown on the year 1 financial statements assuming Munoz used the labor and materials to make 1,500 chairs. Further, assume that Munoz sold 1,200 of the chairs it made.

B) Determine the total amount of assets and the amount of expense shown on the year 1 financial statements assuming Munoz used the labor and materials to provide dental cleaning services to 500 patients.
Answer #1

A) Cost of Goods Produced = $157500+48300 = $205800
Cost of Goods Sold = $205800/1500 x 1200 = $164640
Inventory = $205800-164640 = $41160

Total Amount of Assets = Cash + Inventory + Accounts receivable
= (210000-205800) + 41160 +
= $45360

(Figure of sales not given, so accounts receivable or cash collected on sales ignored)

Total Amount of expense = Cost of Goods sold = $164640

B)  
Total Amount of Assets = Cash + Accounts receivable
= 210000-205800 +
= $4200

(Figure of revenue not given, so accounts receivable or cash collected on sales ignored)

Total Expenses = $205800